OpenStudy (anonymous):

In a group of 212 people, there were 32 more males than females. How many males were in the group?

OpenStudy (anonymous):

Is 122 males correct?

OpenStudy (anonymous):

yeup :)

OpenStudy (butterflydreamer):

Total = 212 Let m = males and f = females There were 32 MORE males than females Males (m) = 32 + f this means Total = (32 + f ) + f 212 = 32 + f + f Now solve for "f" Then plug it into m = 32 + f to find number of males.

OpenStudy (anonymous):

No, need for me to retype the solution!

OpenStudy (anonymous):

@MathNinja_C what?

OpenStudy (anonymous):

Butterflydreamer, type how I got 122 males

OpenStudy (anonymous):

typed*

OpenStudy (anonymous):

oh lol well i got 122 too haha

OpenStudy (butterflydreamer):

excellent ^_^